# Verifying Equilibria in Finite-Horizon Probabilistic Concurrent Game Systems

### v5

@@ 3.4. Strategy Improvements
By definition, we know that if $\pi$ is not a subgame perfect equilibrium, then there exists some history $h$ such that Agent $i$ has a profitable deviation from $\pi|_{h}$ in $\mathbb{G}|_{h}$. We do not know, however, what the strategy that realizes this deviation looks like and if it can be represented in polynomial space. In this section, we address these questions. We also introduce the notation of the value of a state to refer to the value of its hitting probability (see Lemma 3 above), following common texts in Markov Chains such as [Put94].
@@
We now work by backwards induction. Consider the subset of $R$ given by $\langle v,s,n|n=F\rangle$. Clearly, all choices of actions, including the one suggested by $\pi_{i}$, are optimal as these states have no outgoing transitions. Now consider the subset of $R$ given by $\langle v,s,n|n=F-1\rangle$. Since the value at each of the states in $\langle v,s,n|n=F\rangle$ is optimal by default, it can be determined if $\pi_{i}$ suggests an optimal distribution by maximizing over all actions in $A_{i}$$\mathbb{G}_{i}\times\pi$ is an MDP, so at least one of the distributions that achieve an optimal value is deterministic [Put94]. If all such values are indeed optimal, then we continue to the subset of $R$ given by $\langle v,s,n|n=F-2\rangle$. Eventually, we must reach a state $v$ where the value is not optimal. Let $v^{*}=\langle v,s,n^{*}\rangle$ be the first such state observed; if there are multiple states with the same largest time index $n^{*}$, then we sort them lexicographically. We are not interested in solving $\mathbb{G}_{i}\times\pi$ explicitly; this process terminates after a single change has been made at one state $v^{*}$.
@@ 3.6.1. Motivation Behind Deterministic Transition Functions in Strategy Transducers
In Definition 2, we introduced finite-state transducers with deterministic transition functions as models of strategies. In this section, we elaborate on the motivation behind this choice, which boils down to two main points – these transducers simplify our presentation and are mathematically sufficient models. Using a deterministic transition function, we guaranteed that after reading a history $h\in V^{+}$, a transducer $\pi_{i}$ would be a single state $s\in S_{i}$. This allowed us to directly connect subgames to individual states in $\mathbb{G}\times\pi$, simplifying the presentation of Section 3. We also note that it can be shown that Nash equilibria and subgame perfect equilibria always exist in our concurrent game model through the use of backward induction. Backwards induction would result in Markovian strategies for each agent, i.e. strategies of type $V\times[F]\rightarrow\mathbb{D}(A_{i})$ [MT01]. Therefore, considering transducers with deterministic transitions that model functions of type $V^{+}\rightarrow\mathbb{D}(A_{i})$ subsumes the family of Markovian strategies (the value corresponding to $[F]$ can be obtained from the length of an element $h\in V^{+}$) and is therefore sufficient – there is no need to consider more general models, since both equilibria always exist when the model presented is considered.
@@ 4.1. Upper Bound
The next step is to compute the probability of the payoff achieved by the most profitable deviation from Agent $i$. This can be done by explicitly constructing and solving the MDP $\mathbb{G}_{i}\times\pi$ described in Section 3.4 (a step we expicitly avoided in Section 3). In every Markov Decision process, there is a policy (a deterministic, memoryless strategy) that achieves the optimal value [Put94] from every state simultaneously. We can solve for this policy through backwards induction [Put94], and since this solution is guaranteed to return deterministic distributions at each state, there is no issue of representation when it comes to the numerical values used in the policy. Once this optimal policy is fixed, applying Lemma 6 gives us an exponential bound (in the size of the original input of $\mathbb{G}$ and $\pi$) on the size of the representation of the optimal hitting probability from the initial state.
Explicitly represented MDPs can be solved in polynomial time [PT87, Put94]. Therefore, scaling these polynomial time algorithms up to the explicitly constructed exponentially large $\mathbb{G}_{i}\times\pi$ gives us an EXPTIME upper bound for computing the optimal hitting probability from the initial state. Once both hitting probabilities have been computed – the original and the computed optimal, the only remaining step is to compare them. This can be done in EXPTIME since the numbers involved have at most exponential length w.r.t the input. If the optimal value from the initial state in $\mathbb{G}_{i}\times\pi$ is larger than the value from the initial state in $\mathbb{G}\times\pi$, then Agent $i$ has a profitable deviation, and $\pi$ is not a Nash equilibrium. By checking each agent in this way, we can decide whether some agent has a profitable deviation in EXPTIME and, therefore, decide the Nash equilibrium verification problem in EXPTIME.
Broadly speaking, an EXPTIME upper bound for the verification problem appears in various other works in various other forms as a “folk theorem,” e.g. [YSD+22]. In this work, we make the details explicit and consider issues of numerical representation that are often ignored in other works.
@@ 4.2. Lower Bound
The presence of universal states means that the computations in an alternating Turing machine are represented as trees instead of paths. Since universal states require that both successors accept, such transitions are precisely the areas where the tree branches. For more precise details, we refer the reader to [CS76].
@@ 4.2.1. A High Level View of the Reduction
This problem is hard for the class APSPACE = EXPTIME [CS76]. In this decision problem, the machine $M$ can never use more than $n$ cells in an accepting computation, as any computation that tries to access more than that amount of space is immediately rejected. Since the machine can never use more than $n$ cells, we introduce the concept of an ID – a string that represents the complete information of the tape and machine at a given moment in time. {exa}For example, an ID may read $122\langle r_{3},5\rangle 68$. This string conveys the following information : (1) The string is $6$ characters long – $\langle r_{3},5\rangle$ is considered to be a single character. This means that the bounded Turing machine can only use $6$ cells, i.e., $n=6$. (2) The first cell, counting from the left, reads $1^{\prime}$. For this example, $\Gamma$ is given by the set of digits. (3) The head of the Turing machine is currently reading the fourth cell from the left, represented by the character $\langle r_{3},5\rangle$. The machine is in state $r_{3}\in R$, reading the character $^{\prime}5^{\prime}$ written on the fourth cell from the left.
@@
The high-level idea behind the simulation of $M$ is to track its computations by tracking its IDs. The number of IDs of $M$ is, however, exponential in the size of $M$. Therefore, we opt to store them in a distributed manner. The verification problem we construct contains $n$ agents that correspond to the $n$ characters of $M$’s ID strings. The $i$-th character of the ID is stored in the state of $\pi_{i}$. These $n$ agents have a reachability goal of the entire set $V$, the set of states of the game $\mathbb{G}$. Therefore, they always receive a constant payoff of $1$ regardless of what strategies are employed and have no incentive to deviate from the strategies assigned to them by $\pi$. As a shorthand, we refer to these agents as “simulators”, since their only purpose is to simulate the computations of $M$ by tracking its IDs.
@@ 4.2.3. Input Strategy Profile Construction
In this section, we provide details on how to construct the strategy profile $\pi$. Recall that the set of actions for Agents $1$ through $n$ is given by $\{[n]\times\Gamma\times\{\rightarrow,\leftarrow\}\times R\}\cup\{*\}\cup\{X\}$ as defined in Section 4.2.2. For Agent $n+1$, the set of actions is given by $\{\alpha,\beta\}$.
@@ 5. Concluding Remarks and Future Work
If we consider a deviating agent $j$, we can see the fundamental difference between subgame perfect equilibrium and Nash equilibrium – by examining the Markov Decision Process chain $\mathbb{G}_{i}\times\pi$ (see Section 3.4), we can see that the conditions of a strategic improvement for agent $j$ are different. For subgame perfect equilibrium, if the value at some history-relevant state can be improved, then $\pi_{j}$ can be deviated from, and $\pi$ is not a subgame perfect equilibrium. This closely corresponds to the concept of an “optimal policy” from the literature on Markov Decision Processes [Put94], which is a memoryless strategy that maximizes the value of all states simultaneously. For Nash equilibrium, however, the agent is only incentivized to deviate if the value of the initial state specifically can be improved, which does not directly correspond to notions of optimality in the Markov-Decision-Process literature. This observation paints an interesting picture of the relationship between these equilibrium concepts and other strategy-related solution concepts.

### v6

@@ 3.4. Strategy Improvements
By definition, we know that if $\pi$ is not a subgame perfect equilibrium, then there exists some history $h$ such that Agent $i$ has a profitable deviation from $\pi|_{h}$ in $\mathbb{G}|_{h}$. We do not know, however, what the strategy that realizes this deviation looks like and if it can be represented in polynomial space. In this section, we address these questions. We also introduce the notation of the value of a state to refer to the value of its hitting probability (see Lemma 3 above), following common texts in Markov Chains such as [34].
@@
We now work by backwards induction. Consider the subset of $R$ given by $\langle v,s,n|n=F\rangle$. Clearly, all choices of actions, including the one suggested by $\pi_{i}$, are optimal as these states have no outgoing transitions. Now consider the subset of $R$ given by $\langle v,s,n|n=F-1\rangle$. Since the value at each of the states in $\langle v,s,n|n=F\rangle$ is optimal by default, it can be determined if $\pi_{i}$ suggests an optimal distribution by maximizing over all actions in $A_{i}$$\mathbb{G}_{i}\times\pi$ is an MDP, so at least one of the distributions that achieve an optimal value is deterministic [34]. If all such values are indeed optimal, then we continue to the subset of $R$ given by $\langle v,s,n|n=F-2\rangle$. Eventually, we must reach a state $v$ where the value is not optimal. Let $v^{*}=\langle v,s,n^{*}\rangle$ be the first such state observed; if there are multiple states with the same largest time index $n^{*}$, then we sort them lexicographically. We are not interested in solving $\mathbb{G}_{i}\times\pi$ explicitly; this process terminates after a single change has been made at one state $v^{*}$.
@@ 3.6.1. Motivation Behind Deterministic Transition Functions in Strategy Transducers
In Definition 2, we introduced finite-state transducers with deterministic transition functions as models of strategies. In this section, we elaborate on the motivation behind this choice, which boils down to two main points – these transducers simplify our presentation and are mathematically sufficient models. Using a deterministic transition function, we guaranteed that after reading a history $h\in V^{+}$, a transducer $\pi_{i}$ would be a single state $s\in S_{i}$. This allowed us to directly connect subgames to individual states in $\mathbb{G}\times\pi$, simplifying the presentation of Section 3. We also note that it can be shown that Nash equilibria and subgame perfect equilibria always exist in our concurrent game model through the use of backward induction. Backwards induction would result in Markovian strategies for each agent, i.e. strategies of type $V\times[F]\rightarrow\mathbb{D}(A_{i})$ [26]. Therefore, considering transducers with deterministic transitions that model functions of type $V^{+}\rightarrow\mathbb{D}(A_{i})$ subsumes the family of Markovian strategies (the value corresponding to $[F]$ can be obtained from the length of an element $h\in V^{+}$) and is therefore sufficient – there is no need to consider more general models, since both equilibria always exist when the model presented is considered.
@@ 4.1. Upper Bound
The next step is to compute the probability of the payoff achieved by the most profitable deviation from Agent $i$. This can be done by explicitly constructing and solving the MDP $\mathbb{G}_{i}\times\pi$ described in Section 3.4 (a step we expicitly avoided in Section 3). In every Markov Decision process, there is a policy (a deterministic, memoryless strategy) that achieves the optimal value [34] from every state simultaneously. We can solve for this policy through backwards induction [34], and since this solution is guaranteed to return deterministic distributions at each state, there is no issue of representation when it comes to the numerical values used in the policy. Once this optimal policy is fixed, applying Lemma 6 gives us an exponential bound (in the size of the original input of $\mathbb{G}$ and $\pi$) on the size of the representation of the optimal hitting probability from the initial state.
Explicitly represented MDPs can be solved in polynomial time [31, 34]. Therefore, scaling these polynomial time algorithms up to the explicitly constructed exponentially large $\mathbb{G}_{i}\times\pi$ gives us an EXPTIME upper bound for computing the optimal hitting probability from the initial state. Once both hitting probabilities have been computed – the original and the computed optimal, the only remaining step is to compare them. This can be done in EXPTIME since the numbers involved have at most exponential length w.r.t the input. If the optimal value from the initial state in $\mathbb{G}_{i}\times\pi$ is larger than the value from the initial state in $\mathbb{G}\times\pi$, then Agent $i$ has a profitable deviation, and $\pi$ is not a Nash equilibrium. By checking each agent in this way, we can decide whether some agent has a profitable deviation in EXPTIME and, therefore, decide the Nash equilibrium verification problem in EXPTIME.
Broadly speaking, an EXPTIME upper bound for the verification problem appears in various other works in various other forms as a “folk theorem,” e.g. [46]. In this work, we make the details explicit and consider issues of numerical representation that are often ignored in other works.
@@ 4.2. Lower Bound
The presence of universal states means that the computations in an alternating Turing machine are represented as trees instead of paths. Since universal states require that both successors accept, such transitions are precisely the areas where the tree branches. For more precise details, we refer the reader to [10].
@@ 4.2.1. A High Level View of the Reduction
This problem is hard for the class APSPACE = EXPTIME [10]. In this decision problem, the machine $M$ can never use more than $n$ cells in an accepting computation, as any computation that tries to access more than that amount of space is immediately rejected. Since the machine can never use more than $n$ cells, we introduce the concept of an ID – a string that represents the complete information of the tape and machine at a given moment in time. {exa}For example, an ID may read $122\langle r_{3},5\rangle 68$. This string conveys the following information : (1) The string is $6$ characters long – $\langle r_{3},5\rangle$ is considered to be a single character. This means that the bounded Turing machine can only use $6$ cells, i.e., $n=6$. (2) The first cell, counting from the left, reads $1^{\prime}$. For this example, $\Gamma$ is given by the set of digits. (3) The head of the Turing machine is currently reading the fourth cell from the left, represented by the character $\langle r_{3},5\rangle$. The machine is in state $r_{3}\in R$, reading the character $^{\prime}5^{\prime}$ written on the fourth cell from the left.
@@
The high-level idea behind the simulation of $M$ is to track its computations by tracking its IDs. The number of IDs of $M$ is, however, exponential in the size of $M$. Therefore, we opt to store them in a distributed manner. The verification problem we construct contains $n$ agents that correspond to the $n$ characters of $M$’s ID strings. The $i$-th character of the ID is stored in the state of $\pi_{i}$. These $n$ agents have a reachability goal of the entire set $V$, the set of states of the game $\mathbb{G}$. Therefore, they always receive a constant payoff of $1$ regardless of what strategies are employed and have no incentive to deviate from the strategies assigned to them by $\pi$. As shorthand, we refer to these agents as “simulators”, since their only purpose is to simulate the computations of $M$ by tracking its IDs. The general idea behind the representation of the IDs of $M$ is depicted in Figure 4.
@@ 4.2.3. Input Strategy Profile Construction
In this section, we provide details on how to construct the strategy profile $\pi$. Recall that the set of actions for Agents $1$ through $n$ is given by $\{[n]\times\Gamma\times\{\rightarrow,\leftarrow\}\times R\}\cup\{*\}\cup\{X\}$ as defined in Section 4.2.2. For Agent $n+1$, the set of actions is given by $\{\alpha,\beta\}$. Before continuing with the technical details, we refer the reader to Figure 4, which provides a high-level overview of the interaction between $\pi$ and $\mathbb{G}$ in a graphical manner.
Figure 4. A graphical depiction of the interactions between $\pi$ and $\mathbb{G}$. As mentioned before, we use the state space of the individual $\pi_{i}$ to track the ID of $M$. (1) The machine $M$ has some ID, depicted by the top row of small cells. The head is currently over cell $i$, which is in blue. Cells i1i-1 and $i+1$ are depicted in green. The state of $\mathbb{G}$ is $\langle i,\emptyset\rangle\in V_{1}$, indicating that it is Agent $i$’s turn to select an action. To do so, Agent $i+1$ consults the state of his assigned strategy $\pi_{i+1}$. This state records the contents of the cells i1,ii-1,i and $i+1$, noting that they read $a$, $b$, and $c$ in that order. The head is currently over cell $i$ in state $r$. (2) The transition information that is output by $\pi_{i}$ depends on the state. If $r$ is deterministic, the transition information is output in a straightforward manner. If it is existential, then Agent $n+1$ provides input indicating which successor he would like (more information on how this is managed is available in Section 4.2.3). If it is alternating, both successors are output with 50% probability. (3) In this example, the transition information $\langle i,x,\rightarrow,r_{1}\rangle$ is output by $\pi_{i}$. The game $\mathbb{G}$ then moves to the state $\langle i,x,\rightarrow,r_{1}\rangle\in V_{2}$, which allows all agents to see how the ID has updated. In this case, the character $x$ was written at cell $i$, and the head moved right to cell $i+1$ and transitioned to state $r_{1}$. We highlight $\pi_{i},\pi_{i+1}$ and $\pi_{i+2}$, as these are the strategies of the Agents that monitor cell $i+1$ and can therefore see the new position of the head at cell $i+1$. (4) We now take a look at the state of $\pi_{i+1}$ since $i+1$ is the new position of the head. This state has been updated in response to the transition information previously provided by Agent $i$. It is now Agent $i+1$’s turn to output the new transition information, essentially starting back at Step 1.
@@ 5. Concluding Remarks and Future Work
If we consider a deviating agent $j$, we can see the fundamental difference between subgame perfect equilibrium and Nash equilibrium – by examining the Markov Decision Process $\mathbb{G}_{i}\times\pi$ (see Section 3.4), we can see that the conditions of a strategic improvement for agent $j$ are different. For the subgame perfect equilibrium, if the value at some history-relevant state can be improved, then $\pi_{j}$ can be deviated from, and $\pi$ is not a subgame perfect equilibrium. This closely corresponds to the concept of an “optimal policy” from the literature on Markov Decision Processes [34], which is a memoryless strategy that maximizes the value of all states simultaneously. For the Nash equilibrium, however, the agent is only incentivized to deviate if the value of the initial state specifically can be improved, which does not directly correspond to notions of optimality in the Markov-Decision-Process literature. This observation paints an interesting picture of the relationship between these equilibrium concepts and other strategy-related solution concepts.